# Time Transfer: On Optimal Learning Rate and Batch Size In The Infinite Data Limit

### v1

@@ A.7.3 $d_{\mathrm{model}}^{\mathrm{base}}=1024,~{}d_{\mathrm{model}}=256$
A.7.4 $d_{\mathrm{model}}^{\mathrm{base}}=256,~{}d_{\mathrm{model}}=256$
@@
A.7.5 $d_{\mathrm{model}}^{\mathrm{base}}=256,~{}d_{\mathrm{model}}=512$
@@
A.7.6 $d_{\mathrm{model}}^{\mathrm{base}}=256,~{}d_{\mathrm{model}}=1024$
@@
A.8 $\mu$P-averaged optimal learning rate and batch size joint scaling for $d_{\mathrm{model}}^{\mathrm{base}}=256$
@@
A.9 Per-model optimal learning rate and batch size joint scaling
A.9.1 $d_{\mathrm{model}}^{\mathrm{base}}=256$
@@
A.9.2 $d_{\mathrm{model}}^{\mathrm{base}}=1024$
@@
A.10 Per-model validation loss evolution in time depending on batch size with optimally-tuned learning rate
A.10.1 $d_{\mathrm{model}}^{\mathrm{base}}=256$
@@
A.10.2 $d_{\mathrm{model}}^{\mathrm{base}}=1024$
@@
A.11 Learning rate sensitivity in the $\mu$P width limit

### v2

@@ A.8.3 $d_{\mathrm{model}}^{\mathrm{base}}=1024,~{}d_{\mathrm{model}}=256$
A.8.4 $d_{\mathrm{model}}^{\mathrm{base}}=256,~{}d_{\mathrm{model}}=256$
@@
A.8.5 $d_{\mathrm{model}}^{\mathrm{base}}=256,~{}d_{\mathrm{model}}=512$
@@
A.8.6 $d_{\mathrm{model}}^{\mathrm{base}}=256,~{}d_{\mathrm{model}}=1024$
@@
A.9 Fig. 1 with the full set of batch size and token budget values
Figure 15: Same as Fig. 1 with the full set of batch size (left) and token budget (right) values.
A.10 $\mu$P-averaged optimal learning rate and batch size joint scaling for $d_{\mathrm{model}}^{\mathrm{base}}=256$
@@
A.11 Per-model optimal learning rate and batch size joint scaling
A.11.1 $d_{\mathrm{model}}^{\mathrm{base}}=256$
@@
A.11.2 $d_{\mathrm{model}}^{\mathrm{base}}=1024$
@@
A.12 Per-model validation loss evolution in time depending on batch size with optimally-tuned learning rate
A.12.1 $d_{\mathrm{model}}^{\mathrm{base}}=256$
@@
A.12.2 $d_{\mathrm{model}}^{\mathrm{base}}=1024$
@@
A.13 Learning rate sensitivity in the $\mu$P width limit